williammill:

Over the past ten years, the town's population doubled in size. The population is currently 12,000. What was the population ten years ago? Which of the following represents the given word problem if "P" represents the population ten years ago?

Gdeinward:

you must provide which of the following. I would imagine a likely answer could be P*2=12000.

Gdeinward:

And its n the wrong subject. Should be mathematics.

Gdeinward:

or it could be P= 12000/2 it depends on the answer choices that we don't have

AngeI:

Its just asking what the population was 10 years ago, if you used the first equation you would still have to divide the current population by 2. So just do 12000/2=6000 lol